The horizontal and vertical components of a momentum are 25 and 25√3Ns . Find magnitude and direction of resultant momentum vector

Grade:11

2 Answers

Phanindra
6 years ago
P1 = 25 and P2 = 25 × 3^1/2....Pnet = (P1^2 + P2^2)^1/2...Parellelogram law...So you will get the answer as....P(net) = 50 Ns...
Khimraj
6 years ago
P1 = 25 Ns
P2 = 25\sqrt{3} Ns
P = \sqrt{P_{1}^{2} + P_{2}^{2}} = 50 Ns
tan\Theta = P2/P1 = \sqrt{3}
So \Theta = 60o.
Hope it clears.